OpenStudy (anonymous):

what is 9.5% of 616?

OpenStudy (anonymous):

58.52

OpenStudy (anonymous):

y the heck don't u use calculator ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The one I have is super weird

OpenStudy (angela210793):

9.5% of 616? Roseanne pay attention on the solution!!!!!!!! 9.5/100*616=58.52

OpenStudy (anonymous):

do you always use that formula?

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

so in this case 275 iis 12.5 % of what # you would use: 12.5/100* 275?

OpenStudy (angela210793):

12.5/100*x=275 do u know how to solve it now?

OpenStudy (anonymous):

x??????????????

OpenStudy (angela210793):

yea solve for x

OpenStudy (anonymous):

so divivde 12.5 into 275 then multiply by 100

OpenStudy (angela210793):

no 12.5/100*x=275 multiply both sides with 100 12.5x=27500 divide both sides with 12.5 x=27500/12.5 x=2200